A box is sitting on a ramp

angled at 20°. If the box weighs

50 N, what is the normal force

on the box?


1
Expert's answer
2021-11-11T17:07:55-0500

The weight is "P = 10N", angle "\alpha =20\degree", normal force is "N" and friction force is "F_{fr}". Using the second Newton's law in a projection on y-axis, obtain:


"N = P\cos\alpha\\
N = 50N\cdot \cos20\degree \approx 47N"

Answer. 47N.